OpenStudy (anonymous):

The point A(-8, 6) is translated using T: (x,y) → (x + 5. y - 4). What is the distance from A to A'? a. 41 b. 21 c. 3 d. 29

OpenStudy (anonymous):

Use distance formula: \[\sqrt{(x _{2} - x _{1} ) + (y _{2} - y _{1} )}\] \[\sqrt{(-3+8)^2 + (2-6)^2}\] \[\sqrt{(5)^2 +(-4)^2}\] \[\sqrt{25+16}\] \[\sqrt{41}\] a is the answer